If an amount of 2,160, which includes 10% simple interest for 2 years, is paid back, how much was borrowed 2 years earlier?( $1800)

Answer by math_helper(2461)

You got it, $1800.

The work-out:
x = amount invested 2yrs ago
2(0.10)x + x = 2160
1.20x = 2160
x = 2160/1.2 = 1800